# Spins of Black Holes in X-ray Binaries and the Tension with the Gravitational Wave Measurements

### v3

@@ 4.1 Revised spin measurements and improved disk models
In the case of GX 339–4, its mass and distance are only roughly constrained (Heida et al., 2017; Zdziarski et al., 2019). Still, the continuum method, coupled with the reflection, has been applied to a study of this source by Parker et al. (2016) and Zdziarski et al. (2025). In the latter, the joint fitting of two sets of high-quality data in soft states with only weak high-energy tails from NICER and NuSTAR has allowed the authors to determine the BH mass and spin and the source distance. However, it turns out that the results strongly depend on the disk model used. The models treating departures from the local blackbodies by a color correction, in particular kerrbb and kerrbb2, yielded strongly negative spins and low masses. However, models employing radiative transfer calculations, namely bhspec and slimbh, yielded much better fits with moderately positive spins and relatively high masses, in agreement with the mass function of Heida et al. (2017). The good fits of these two models show that the observed spectra agree much better with the atmospheric models than those employing color corrections. This is despite the standard disk model (assumed in the atmospheric models) disagreeing with the stability of disks in the soft state (Section 3.1.2). In particular, the disk stability is observed in these observations of GX 339–4 (with the rms variability of the disk emission $\lesssim$1%).
The study of Zdziarski et al. (2025) has also shown that even for fixed BH mass, spin, and distance, different available standard models (without including a warm corona) give significantly different spins; see Table 3 in Zdziarski et al. (2025). In that example, the differences between the spins fitted by different disk-continuum codes are up to $\Delta a_{*}\approx 0.4$. Thus, the published results of fitting the standard continuum disk models should be considered significantly uncertain.
@@
On the other hand, adding a warm corona in the soft state of Cyg X-1 reduces the spin dramatically, to $a_{*}\approx 0$–0.1 (Belczyński et al., 2024; Zdziarski et al., 2024b), see Figure 6. For GX 339–4, it yielded the spin only weakly constrained, $0\lesssim a_{*}\lesssim 0.8$, thus consistent with being very low. When adding a warm corona in LMC X-1, the spin becomes weakly constrained (due to the observed flux of this source being much weaker than that of Cyg X-1), with $a_{*}\approx 0.1$–0.9 (Zdziarski et al., 2024a). Similarly, adding it in M33 X-7 leads to an almost unconstrained spin, $0\lesssim a_{*}\lesssim 0.9$ (Belczyński et al., 2024).
@@ 4.2 Spin-up by accretion
To spin up a BH with an initial mass $M_{0}$ from $a_{*}=0$ to $a_{*}=1$, a mass of $M_{\rm accr}=3M_{0}\arccos\sqrt{2/3}\approx 1.846M_{0}$ needs to be accreted (Bardeen, 1970; Thorne, 1974). It is assumed that the matter accreted carries the specific angular momentum of the ISCO, and the decelerating power of jets (Lowell et al., 2024, 2025) is neglected. However, the BH mass grows by a lower amount, $M_{1}-M_{0}=(6^{1/2}-1)M_{0}\approx 1.449M_{0}$, due to the rest energy being converted to the binding energy. Any additional accretion does not increase the spin. In terms of the final BH mass, the accreted fraction needed to spin up to $a_{*}=1$ is $M_{\rm accr}/M_{1}=\sqrt{2/3}\arccos\sqrt{2/3}\approx 0.754$, while $M_{0}/M_{1}=6^{-1/2}\approx 0.408$. The actual maximum spin is lower than 1 due to the radiation swallowed by the BH, $a_{*{\rm max}}\approx 0.998$ (Thorne, 1974). For $M_{1}/M_{0}\leq 6^{1/2}$, a null initial spin and an arbitrary final spin, $a_{*}$, we have (Bardeen, 1970; Thorne, 1974)
@@
where Equation (11) gives $a_{*}$ as a function of $M_{\rm accr}$, and Equation (12) is the inverse function of Equation (10). Figure 7(a) shows the fractional accreted mass needed to spin up from 0 to $a_{*}$ and the corresponding (lower) mass gain by the BH, and Figure 7(b) shows the spin vs. the accreted and gained masses. We see that the required accreted mass sharply decreases for $a_{*}<1$, e.g., $M_{\rm accr}\approx 0.677M_{0}$ and $0.358M_{0}$ for $a_{*}=0.9$ and 0.7, respectively. We see that a substantial mass is still needed to spin up a BH to a high $a_{*}$. If this explains the BH HMXBs’ spins, we would then need to understand why most of the BBHs were not significantly spun up.
@@ 4.2.1 BH HXMBs
where $X$ is the H mass fraction. At $\eta\approx 0.1$, $X\approx 0.7$, it is $\approx$40 Myr. For an increase of the mass corresponding to a spin increase from 0 to 0.9, $M_{1}/M_{0}\approx 1.68$ (see Figure 7 and the text above), the required time is $t_{\rm Edd}\ln(M_{1}/M_{0})\approx 20$ Myr. This is much longer than the typical lifetime of a BH HMXB, e.g., $\approx$4 Myr for Cyg X-1 (Miller-Jones et al., 2021). This conclusion was also achieved by a recent study of Xing et al. (2025), in which the conservative hypercritical accretion was confirmed (including the Roche-lobe overflow accretion channel) to be required to account for the measured spins of the three known BH HMXBs. Thus, we need hypercritical accretion (as it happens only episodically) if the high spins of those systems are real and their natal spins were low.
@@
The need for hypercritical accretion for the BH HMXBs was also pointed out by Moreno Méndez et al. (2008), Moreno Méndez (2011), and Qin et al. (2022). They considered energy loss from the accretion flow via the formation of neutrino pairs, which they claimed causes accretion to be fully conservative ($\dot{M}_{\rm accr}=\dot{M}_{\rm tr}$). This appears surprising; neutrino pair cooling is important in supernovae and GRBs, but the densities in hypercritical accretion flows are much lower, which will cause the flow to be optically thin to that pair production. Indeed, the simulation performed in Zdziarski et al. (2024b) using the peak transfer rate for Cyg X-1 of the model of Qin et al. (2022), $\dot{M}_{\rm out}=10^{-2}{M}_{\odot}$/yr, showed that the neutrino luminosity was only $\sim 10^{-5}\dot{M}_{\rm out}c^{2}$, i.e., completely negligible.
@@
where $k$ is the Boltzmann constant, $m_{\rm e}$ is the electron mass and $h$ is the Planck constant. At $\dot{M}_{\rm out}=10^{-2}{M}_{\odot}$/yr estimated by Qin et al. (2022) for Cyg X-1, the average electron density within the spherization radius is $n_{\rm e}\sim 3\times 10^{14}$ cm-3 (using the formalism of Poutanen et al. 2007), which gives $T_{\rm F}\approx 0.2$ K. Obviously, the actual electron temperature of that flow is $T_{\rm e}\ggg T_{\rm F}$. Consequently, the gas is nondegenerate, whose chemical potential is negative,
@@ 4.2.2 BH LXMBs
The minimum (corresponding to the fully conservative accretion) donor mass at the onset of the mass transfer onto the BH, $M_{2,0,min}$, required to spin-up the BH from null spin to $a_{*}$ at the current binary masses $M_{1}$ and $M_{2}$ equals $M_{\rm accr}(a_{*},M_{1})+M_{2}$. For given $M_{1}$ and $a_{*}$, $M_{0}$ can be obtained by numerically solving Equation (9), which can then be plugged into Equation (12), which gives $M_{\rm accr}(M_{0},M_{1})$. This yields
@@
An example is shown in Figure 8. In it, $a_{*}=0.5$ yields $M_{0}\approx 5.3{M}_{\odot}$ and $M_{\rm 2,0,min}\approx 1.9{M}_{\odot}$. We also point out that spinning up a BH with a moderate mass to a high $a_{*}$ implies a small initial mass, which then can fall in the range of neutron-star masses. For example, the values of $M_{1}=6{M}_{\odot}$ and $a_{*}=0.998$ imply $M_{0}\approx 2.7{M}_{\odot}$.
@@
Binaries with low-mass companions ($M_{2}\sim 1{M}_{\odot}$) may be more easily disrupted during BH formation, for example, due to the Blaauw kick associated with sudden mass loss (Blaauw, 1961). This effect could reduce the survival probability of such systems and suppress their contribution to the observed BH LMXB population, thus favoring intermediate-mass companions as the progenitors of present-day BH LMXBs. On the other hand, another population of systems hosting BH candidates–including all three detached Gaia BH systems announced so far (El-Badry et al., 2023b; Chakrabarti et al., 2023; El-Badry et al., 2023a; Gaia Collaboration et al., 2024)–contains companions with masses of the order of $\sim 1{M}_{\odot}$. These systems are found in the Galactic field and show no evidence for triple or higher-order multiplicity; their properties also pose a significant challenge for formation through conventional mass-transfer scenarios (Nagarajan et al., 2025; Olejak et al., 2025). Consequently, both observations and evolutionary models currently provide only weak constraints on the initial masses of the lower-mass companions in LMXB progenitors.
@@ 4.3 BBHs and BH HMXBs as two different binary populations
In the case of the known BH HMXBs, the future evolution of Cyg X-1 will certainly lead to the formation of a BBH system, but the BHs will not merge in the Hubble time if the accretion is Eddington-limited. Only for fully conservative accretion, the system will merge within $\gtrsim$5 Gyr (Ramachandran et al., 2025). M33 X-7 will, most likely, undergo a common-envelope stage during which the donor and the BH will merge; thus, it will not become a BBH (Ramachandran et al., 2022). While the nature of the compact object in Cyg X-3 is not certain, it is most likely a BH (Antokhin et al., 2022). However, its masses remain relatively uncertain. If the mass of the WR donor is $\gtrsim 13{M}_{\odot}$, it will become a BBH within less than 1 Myr, and the BHs will most likely merge (Belczyński et al., 2013).
The BHs in mergers of field binaries also have undergone evolution in mass and spin, and in this respect are similar to XRBs. The crucial issue, which we stress in the paper, is that the GW results strongly imply that the natal spins are low. To explain the spins measured in HMXBs, the natal spin has to be very high, in contrast to similar BHs undergoing mergers. No viable explanation has been proposed yet. For LMXBs, either the natal spins are high, which again is unexplained, or the donor masses at the onset of the RLO were high, which remains to be shown, and which we discuss in detail.
@@
Finally, as discussed above, BH progenitors in both populations undergo mass-transfer phases that likely remove a substantial fraction of the hydrogen envelope. In addition, XRBs are frequently found in relatively high-metallicity environments, where stellar winds can remove up to $\sim 80\%$ of the progenitor’s initial mass (Belczynski et al., 2010), carrying away a significant amount of angular momentum. If the failed supernova spin-up scenario were operating efficiently, one would naively expect GW sources—often forming in lower-metallicity environments, experiencing weaker stellar winds (and thus more likely retaining part of their hydrogen envelopes)—to host more rapidly spinning BHs than those observed in XRBs. This expectation is contrary to current observations.
@@ 5 Discussion and summary
(i) Natal spins of BHs in merger events are found to be low. As studies of GWs have found, BBHs are characterized by low spins on average. This is especially the case for the first-born BHs; see Figure 1. This also implies that the natal spins of first-born BHs are low.

### v4

@@ 4.1 Revised spin measurements and improved disk models
In the case of GX 339–4, its mass and distance are only roughly constrained (Heida et al., 2017; Zdziarski et al., 2019). Still, the continuum method, coupled with the reflection, has been applied to a study of this source by Parker et al. (2016) and Zdziarski et al. (2025). In the latter, the joint fitting of two sets of high-quality data in soft states with only weak high-energy tails from NICER and NuSTAR has allowed the authors to determine the BH spin. However, it turns out that the results strongly depend on the disk model used. The models treating departures from the local blackbodies by a color correction, in particular kerrbb and kerrbb2, yielded strongly negative spins. However, models employing radiative transfer calculations, namely bhspec and slimbh, yielded much better fits with moderately positive spins, and relatively high masses, in agreement with the mass function of Heida et al. (2017). The good fits of these two models show that the observed spectra agree much better with the atmospheric models than those employing color corrections. This is despite the standard disk model (assumed in the atmospheric models) disagreeing with the stability of disks in the soft state (Section 3.1.2). In particular, the disk stability is observed in these observations of GX 339–4 (with the rms variability of the disk emission $\lesssim$1%).
The study of Zdziarski et al. (2025) has also shown that even for fixed BH mass and distance, different available standard models (without including a warm corona) give significantly different spins; see Table 3 in Zdziarski et al. (2025). In that example, the differences between the spins fitted by different disk-continuum codes are up to $\Delta a_{*}\approx 0.4$. Thus, the published results of fitting the standard continuum disk models should be considered significantly uncertain.
@@
On the other hand, adding a warm corona in the soft state of Cyg X-1 reduces the spin dramatically, to $a_{*}\approx 0$–0.1 (Belczyński et al., 2024; Zdziarski et al., 2024b), see Fig. 6. For GX 339–4, it yielded the spin only weakly constrained, $0\lesssim a_{*}\lesssim 0.8$, thus consistent with being very low. When adding a warm corona in LMC X-1, the spin becomes weakly constrained (due to the observed flux of this source being much weaker than that of Cyg X-1), with $a_{*}\approx 0.1$–0.9 (Zdziarski et al., 2024a). Similarly, adding it in M33 X-7 leads to an almost unconstrained spin, $0\lesssim a_{*}\lesssim 0.9$ (Belczyński et al., 2024).
@@ 4.2 Spin-up by accretion
To spin up a BH with an initial mass $M_{0}$ from $a_{*}=0$ to $a_{*}=1$, a mass of $M_{\rm accr}=3M_{0}\arccos\sqrt{2/3}\approx 1.846M_{0}$ needs to be accreted (Bardeen, 1970; Thorne, 1974). It is assumed that the matter accreted carries the specific angular momentum of the ISCO, and the decelerating power of jets (Lowell et al., 2024, 2025) is neglected. However, the BH mass grows by a lower amount, $M_{1}-M_{0}=(6^{1/2}-1)M_{0}\approx 1.449M_{0}$, due to the rest energy being converted to the binding energy. Any additional accretion does not increase the spin. In terms of the final BH mass, the accreted fraction needed to spin up to $a_{*}=1$ is $M_{\rm accr}/M_{1}=\sqrt{2/3}\arccos\sqrt{2/3}\approx 0.754$, while $M_{0}/M_{1}=6^{-1/2}\approx 0.408$. For $M_{1}/M_{0}\leq 6^{1/2}$, a null initial spin and an arbitrary final spin, $a_{*}$, we have (Bardeen, 1970; Thorne, 1974)
@@
where Equation (11) gives $a_{*}$ as a function of $M_{\rm accr}$, and Equation (12) is the inverse function of Equation (10). Fig. 7(a) shows the fractional accreted mass needed to spin up from 0 to $a_{*}$ and the corresponding (lower) mass gain by the BH, and Fig. 7(b) shows the spin vs. the accreted and gained masses. We see that the required accreted mass sharply decreases for $a_{*}<1$, e.g., $M_{\rm accr}\approx 0.677M_{0}$ and $0.358M_{0}$ for $a_{*}=0.9$ and 0.7, respectively. We see that a substantial mass is still needed to spin up a BH to a high $a_{*}$. If this explains the BH HMXBs’ spins, we would then need to understand why most of the BBHs were not significantly spun up.
@@ 4.2.1 BH HXMBs
where $X$ is the H mass fraction. At $\eta\approx 0.1$, $X\approx 0.7$, it is $\approx$40 Myr. For an increase of the mass corresponding to a spin increase from 0 to 0.9, $M_{1}/M_{0}\approx 1.68$ (see Fig. 7 and the text above), the required time is $t_{\rm Edd}\ln(M_{1}/M_{0})\approx 20$ Myr. This is much longer than the typical lifetime of a BH HMXB, e.g., $\approx$4 Myr for Cyg X-1 (Miller-Jones et al., 2021). This conclusion was also achieved by a recent study of Xing et al. (2025), in which the conservative hypercritical accretion was confirmed (including the Roche-lobe overflow accretion channel) to be required to account for the measured spins of the three known BH HMXBs. Thus, we need hypercritical accretion (as it happens only episodically) if the high spins of those systems are real and their natal spins were low.
@@
The need for hypercritical accretion for the BH HMXBs was also pointed out by Moreno Méndez et al. (2008), Moreno Méndez (2011), and Qin et al. (2022). They considered energy loss from the accretion flow via the formation of neutrino pairs, which they claimed causes accretion to be fully conservative ($\dot{M}_{\rm accr}=\dot{M}_{\rm tr}$). This appears surprising; neutrino pair cooling is important in supernovae and GRBs, but the densities in hypercritical accretion flows are much lower, which will cause the flow to be optically thin to that pair production. Indeed, the simulation performed in Zdziarski et al. (2024b) using the peak transfer rate for Cyg X-1 of the model of Qin et al. (2022), $\dot{M}_{\rm tr}=10^{-2}{M}_{\odot}$/yr, showed that the neutrino luminosity was only $\sim 10^{-5}\dot{M}_{\rm tr}c^{2}$, i.e., completely negligible.
@@
where $k$ is the Boltzmann constant, $m_{\rm e}$ is the electron mass and $h$ is the Planck constant. At $\dot{M}_{\rm tr}=10^{-2}{M}_{\odot}$/yr estimated by Qin et al. (2022) for Cyg X-1, the average electron density within the spherization radius is $n_{\rm e}\sim 3\times 10^{14}$ cm-3 (using the formalism of Poutanen et al. 2007), which gives $T_{\rm F}\approx 0.2$ K. Obviously, the actual electron temperature of that flow is $T_{\rm e}\ggg T_{\rm F}$. Consequently, the gas is nondegenerate, whose chemical potential is negative,
@@ 4.2.2 BH LXMBs
The minimum (corresponding to the fully conservative accretion) donor mass at the onset of the mass transfer onto the BH, $M_{\rm 2,0,min}$, required to spin-up the BH from null spin to $a_{*}$ at the current binary masses $M_{1}$ and $M_{2}$ equals $M_{\rm accr}(a_{*},M_{1})+M_{2}$. For given $M_{1}$ and $a_{*}$, $M_{0}$ can be obtained by numerically solving Equation (9), which can then be plugged into Equation (12), which gives $M_{\rm accr}(M_{0},M_{1})$. This yields
@@
An example is shown in Fig. 8. In it, $a_{*}=0.5$ yields $M_{0}\approx 5.3{M}_{\odot}$ and $M_{\rm 2,0,min}\approx 1.9{M}_{\odot}$. We also point out that spinning up a BH with a moderate mass to a high $a_{*}$ implies a small initial mass, which then can fall in the range of neutron-star masses. For example, the values of $M_{1}=6{M}_{\odot}$ and $a_{*}=0.998$ imply $M_{0}\approx 2.7{M}_{\odot}$.
@@
Binaries with low-mass companions ($M_{2}\sim 1{M}_{\odot}$) may be more easily disrupted during BH formation, for example, due to the Blaauw kick associated with sudden mass loss (Blaauw, 1961). This effect could reduce the survival probability of such systems and suppress their contribution to the observed BH LMXB population, thus favoring intermediate-mass companions as the progenitors of present-day BH LMXBs. On the other hand, another population of systems hosting BHs–including all three detached Gaia BH systems announced so far (El-Badry et al., 2023b; Chakrabarti et al., 2023; El-Badry et al., 2023a; Gaia Collaboration et al., 2024)–contains companions with masses of the order of $\sim 1{M}_{\odot}$. These systems are found in the Galactic field and show no evidence for triple or higher-order multiplicity; their properties also pose a significant challenge for formation through conventional mass-transfer scenarios (Nagarajan et al., 2025; Olejak et al., 2025). Consequently, both observations and evolutionary models currently provide only weak constraints on the initial masses of the lower-mass companions in LMXB progenitors.
@@ 4.3 BBHs and BH HMXBs as two different binary populations
In the case of the known BH HMXBs, the future evolution of Cyg X-1 will certainly lead to the formation of a BBH system, but the BHs will not merge in the Hubble time if the accretion is Eddington-limited. Only for fully conservative accretion, the system will merge within $\gtrsim$5 Gyr (Ramachandran et al., 2025). M33 X-7 will, most likely, undergo a common-envelope stage during which the donor and the BH will merge; thus, it will not become a BBH (Ramachandran et al., 2022). While the nature of the compact object in Cyg X-3 is not certain, it is most likely a BH (Antokhin et al., 2022). However, its masses remain uncertain. If the mass of the WR donor is $\gtrsim 13{M}_{\odot}$, it will become a BBH within less than 1 Myr, and the BHs will most likely merge (Belczyński et al., 2013).
The BHs in mergers of field binaries also have undergone evolution in mass and spin, and in this respect are similar to XRBs. The crucial issue, which we stress in this paper, is that the GW results strongly imply that the natal spins are low. To explain the spins measured in HMXBs, the natal spin has to be very high, in contrast to similar BHs undergoing mergers. No viable explanation has been proposed yet. For LMXBs, either the natal spins are high, which again is unexplained, or the donor masses at the onset of the RLOF were high, which remains to be shown, and which we discuss in detail.
@@
Finally, as discussed above, BH progenitors in both populations undergo mass-transfer phases that likely remove a substantial fraction of the hydrogen envelope. In addition, XRBs are frequently found in relatively high-metallicity environments, where stellar winds can remove up to $\sim 80\%$ of the progenitor’s initial mass (Belczyński et al., 2010), carrying away a significant amount of angular momentum. If the failed supernova spin-up scenario were operating efficiently, one would naively expect GW sources—often forming in lower-metallicity environments, experiencing weaker stellar winds (and thus more likely retaining part of their hydrogen envelopes)—to host more rapidly spinning BHs than those observed in XRBs. This expectation is contrary to current observations.
@@ 5 Discussion and summary
(i) Natal spins of BHs in merger events are found to be low. As studies of GWs have found, BBHs are characterized by low spins on average. This is especially the case for the first-born BHs; see Fig. 1. This also implies that the natal spins of first-born BHs are low.